OpenStudy (anonymous):

R=2/3t+v solve for t? Don't you add v on both sides, leaving you with r+v=2/3t but what do I do with that fraction?

OpenStudy (anonymous):

\[R=\frac{2}{3t+v}\] \[R(3t+v)=\frac{2}{3t+v}(3t+v)=2\] \[3Rt+Rv=2\] \[3Rt=2-Rv\] \[t=\frac{2-Rv}{3R}=\frac{2}{3R}-\frac{v}{3}\]

OpenStudy (anonymous):

just remember when you need to get something out of the denominator, multiply both sides by the denominator, and continue solving

OpenStudy (anonymous):

Sorry like this R=2 / 3t + v